OpenStudy (anonymous):

find the inverse of f(x)=3 squareroot of x+3

OpenStudy (amistre64):

using () to indicate what is under the radical for clarity ... 3 sqrt(x+3) i assume?

OpenStudy (anonymous):

Yes!

OpenStudy (amistre64):

y = 3 sqrt(x+3) ; solve for x really .... divide out the 3 y/3 = sqrt(x+3) ; ^2 the dies to get rid of the sqrt (y/3)^2 = x+3 ; now -3 to get x all alone (y/3)^2 - 3 = x ; and that you inverse

OpenStudy (anonymous):

f-1(x) = (x/3)^2-3

OpenStudy (amistre64):

that almost the inverse; we have to restrict it to match up tho

OpenStudy (anonymous):

I am writing all of this down to see if I understand

OpenStudy (amistre64):

(y/3)^2 - 3 produces two values for every input; which makes it not one to one material

OpenStudy (amistre64):

we have to restrict it to y>=0 by convention

OpenStudy (anonymous):

Wow. I am going over the problem real quick

OpenStudy (amistre64):

y>=-3 maybe better :)

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x+3}\]

OpenStudy (anonymous):

\[f^{-1}(x)=x^3-3\]

OpenStudy (anonymous):

no restrictions need apply

OpenStudy (anonymous):

wow!! Ok, still looking just to make sure I understand

OpenStudy (amistre64):

there is debate over whether the inverse is properly a function of y as opposed to x

OpenStudy (anonymous):

i am assuming it is \[f(x)=\sqrt[3]{x+3}\] not \[f(x)=3\sqrt{x+3}\]

OpenStudy (amistre64):

y = f(x) then its inverse is f-1(y) = x or some such

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x+3}\] has domain all real number and range as well. there is no restriction unless you put one on it for some other reason, like say x is a length and has to be positive

OpenStudy (anonymous):

So is f-1 (x)=x3-3?

OpenStudy (anonymous):

Is this the answer?

OpenStudy (anonymous):

it is \[f^{-1}(x)=x^3-3\] with no restriction on domain or range

OpenStudy (anonymous):

your original function f says add 3 take the cube root your inverse will then say cube subtract 3

OpenStudy (anonymous):

Ok, thank you! Ya'll are so smart! Can someone help me with this one?

OpenStudy (anonymous):

opposite operations in reverse order

OpenStudy (anonymous):

i am assuming it is CUBE ROOT not 3 times SQUARE ROOT

OpenStudy (anonymous):

Find the equation of the line in general form (5,-4) (-4,3) we know up until the slope which we get -7/9 but then what numbers do you input for the formula?

OpenStudy (anonymous):

sensei is right if it is the second one.

OpenStudy (amistre64):

in any case; the inverse is when you rewrite to problem to solve for the other variable

OpenStudy (anonymous):

you ccan use either coordinate for theh formula as long as you put in y, m and x and leave b as a variable to be firgured out

OpenStudy (amistre64):

P(x,y); Px, Py y = mx -mPx +Py

OpenStudy (anonymous):

use the almighty point slope formula \[y-y_1=m(x-x_1)\] with \[x_1=5, y_1=-4, m=-\frac{7}{9}\]

OpenStudy (amistre64):

but standard form is is ax+by where slope = -a/b

OpenStudy (anonymous):

y=mx+b 3=(-7/9)-4+b 3=28/9+b solve for b...

OpenStudy (anonymous):

standard form for me is \[y=mx+b\] but i am surely wrong

OpenStudy (anonymous):

i like slope intersect form...

OpenStudy (amistre64):

7x +9y = n; plug in a point to calibrate for n

OpenStudy (anonymous):

satellite, thats slope intersect... and btw you got my calc problem wrong ;)

OpenStudy (anonymous):

whatever you are in to. you now have 3 essentially the same way of doing it. good luck

OpenStudy (anonymous):

did not!

OpenStudy (anonymous):

did too

OpenStudy (anonymous):

really? where was the mistake?

OpenStudy (amistre64):

(5,-4) 7(4) +9(-4) = n 28 -36 = n -8 = n 7x +9y = -8

OpenStudy (anonymous):

you divided by two somewhere...

OpenStudy (anonymous):

i am fairly sure the volume was \[V(x)=x(7-2x)(10-2x)\] eys?

OpenStudy (anonymous):

your answer gave me 1.3blah blah whereas the real answer is 2.7blahblah

OpenStudy (anonymous):

you got the equations right, you solved for 0 wrong lol

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Wow! U guys r too smart for me to even figure out why you are arguing

OpenStudy (anonymous):

\[V(x)=4 x^3-34 x^2+70 x\] yes?

OpenStudy (anonymous):

\[V'(x)=12x^2-64x+70\] for sure

OpenStudy (anonymous):

Someone else said that 7x+9y=-1

OpenStudy (anonymous):

now i set = 0 and solve and get oh oops

OpenStudy (anonymous):

But someone here said it is -8

OpenStudy (anonymous):

lol see you got it wrong :P